# Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young M4 + [T7 + T8] hierarchical triple

### v1

@@ Abstract
Benchmark brown dwarfs in wide binary systems are crucial for characterizing substellar objects and calibrating atmospheric and evolutionary models. However, brown dwarf benchmarks with subsolar metallicity, very cool temperatures, or suitability for dynamical mass measurements are rare, limiting our understanding across the full range of mass, age, and metallicity. We present the discovery of two new multiple systems containing T dwarf companions, identified through a targeted search using CatWISE and Gaia databases. L 122-88 AB is a wide binary comprising a mildly metal-poor M2 dwarf and a T5 dwarf, separated by 215.6 arcsec at a distance of 33.1060.014 pc. Atmospheric model fitting to the near infrared spectrum of L 122-88 A suggests a mildly metal-poor composition ([Fe/H] = 0.2). UPM J10403551 AB is a candidate hierarchical triple system at 25.2830.013 pc, consisting of an M4 dwarf and a probable unresolved spectral binary of T7 and T8 dwarfs, separated by 65.48 arcsec from the primary. The H emission detected in UPM J10403551 A indicates an age range of 0.3-2.0 Gyr. This age estimate suggests that the T8 component has a mass between 9 and 28 Jupiter masses, potentially classifying it as a planetary-mass object. These systems augment the sample of benchmark brown dwarfs, particularly in the underexplored regime of cool temperature, providing valuable opportunities for refining our understanding of substellar objects.
@@
binaries: general – brown dwarfs – stars: late-type – stars: low-mass – gaseous planets1 Introduction
@@
Brown dwarfs (BDs) are the most recently discovered major population on the Hertzsprung-Russell (H-R) diagram. They mostly form like stars but lack sufficient mass to sustain steady hydrogen fusion. Consequently, BDs represent the low-mass tail of the initial mass function (IMF) and are crucial for constraining it (Kirkpatrick et al., 2024). Bridging the gap between stars and planets, BDs share similar physical and atmospheric properties with gas giant planets (Hatzes & Rauer, 2015), providing essential references for the characterization of exoplanets.
@@
To address the scientific objectives outlined above, samples of well-characterized BDs are required. However, BDs are challenging to characterize due to their faintness, cool temperatures, and the mass–age degeneracy inherent in their evolution. Without steady hydrogen burning, BDs cool over time, altering their spectral features and types (e.g. fig. 8 in Burrows et al., 2001). Therefore, benchmark BD samples with known mass, age, metallicity, and luminosity, spanning a wide parameter space, are needed to test atmospheric and evolutionary models. This need is especially pressing in the era of ongoing and upcoming deep surveys such as Euclid (Laureijs et al., 2011), the Roman Space Telescope (Spergel et al., 2015), the Rubin Observatory (LSST Science Collaboration et al., 2017), and the Chinese Space Station Telescope (CSST, Zhan, 2011).
@@
BDs in nearby open clusters or moving groups with well-established ages serve as valuable age benchmarks (e.g., Teide 1, Rebolo et al. 1995; Luhman 16 AB, Luhman 2013). However, these clusters and groups are predominantly young ( 1 Gyr) and relatively distant ( 100 pc). In contrast, field BDs constitute the majority of the BD population, span a broader range of ages and metallicities, with many residing in the solar neighbourhood. Yet, direct measurements of age or mass for individual field BDs remain elusive. Nonetheless, BDs that are wide companions to well-characterized stars provide ideal benchmarks (e.g., Gliese 229 B, Nakajima et al. 1995; Indi B, Scholz et al. 2003; McCaughrean et al. 2004; King et al. 2010). Binary systems are expected to have coeval formation, sharing common age, metallicity, and distance—parameters more readily determined for stars than for BDs. This characteristic makes such systems particularly valuable for BD studies.
@@
The T dwarf spectral class was established to classify BDs cool enough to exhibit methane absorption in their near-infrared (NIR) spectra (Burgasser et al., 2002; Geballe et al., 2002). T dwarfs have effective temperatures () between 500–1300 K (Kirkpatrick et al., 2021). They constitute the majority of field BDs (e.g. Zhang et al., 2019). However, T dwarfs are relatively rare in observations due to their cool temperatures and low luminosities. About 900 spectroscopically confirmed T dwarfs are known to date, mostly within 30 pc of the Sun (see summary in Best et al., 2024). Two-thirds of known T dwarfs have spectral types of T5–T9, owing to the rapid evolution of BDs through the L/T transition (e.g. Best et al., 2021).
@@
Currently, there are about 30 T dwarfs known with stellar companions (summarised in Best et al., 2024), with significant contributions from the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010). To expand the parameter space of benchmark T dwarfs and improve their characterization, we searched for late-type T dwarfs with wide stellar companions using the CatWISE2020 catalog (Marocco et al., 2021) and the third data release (DR3) of Gaia (Gaia Collaboration et al., 2023). In this paper, we present the discovery of two new M + T dwarf wide binary systems. Section 2 describes the selection of wide binary candidates. Spectroscopic observation and characterization of these two systems are presented in Sections 3 and 4, respectively. A summary and conclusions are presented in Section 5.
@@ 2 Selection of wide binaries
Wide binaries ( 100 au) exhibit negligible orbital motion over decadal timescales due to their extremely long orbital periods ( 1 kyr). In the solar neighbourhood ( 100 pc), the proper motions (PMs) of wide binary components significantly exceed their orbital motions, resulting in apparent common PM over decades. This characteristic makes common PM an effective method for identifying nearby wide binaries (e.g., Zhang, 2019; Zhang et al., 2024).
T dwarfs emitting primarily in the infrared, only those in the solar neighbourhood are detectable by large-scale infrared surveys. The CatWISE database, providing mid-infrared photometry and PMs for nearby objects, serves as a powerful tool for identifying nearby T dwarf wide binaries.
We focused on late-type T dwarfs, which exhibit redder colour (Eq. 1) and suffer less contamination, e.g. from poorly measured background objects, than earlier types. Candidate selection from CatWISE employed the following criteria:
In our selection, we required signal to noise ratio (SNR) of better than 5.0 for W1 and W2 detections (Eq. 2). SNR of 5.0 corresponds to an error of 0.22 in W1 or W2 magnitudes. To mitigate contamination from background stars in the low-resolution (275 per pixel) unWISE images (Lang, 2014) used for CatWISE, we excluded the Galactic plane (Eq. 5). We select objects without identified artifacts in both W1 and W2 band unWISE images (Eq. 7) and with only one blend component used in the fit (Eq. 6). Approximately 9,540 objects in CatWISE met these criteria (Eqs 1–7).
@@
Given the relatively large PM errors in CatWISE, we required significant PM (Eq. 3) but relaxed the PM constraint in the GCNS cross-match (Eq. 8). The distance limit of 40 pc (Eq. 9) corresponds to the -band detection limit of CatWISE for late T dwarfs (e.g. table 6 of Zhang et al., 2019). We limited the binary separation to within 10′ (corresponding to 24,000 au at 40 pc), beyond which the number of random contaminations arises, partially due to the large PM errors in CatWISE.
@@
Our search yielded 15 candidate common PM pairs. After visual inspection using data from the Dark Energy Survey (DES; Abbott et al., 2021) and the DESI Legacy Imaging Surveys (Dey et al., 2019a), we excluded seven pairs of which the faint component has poor unWISE images, has Dark Energy Camera (DECam) deep -band non-detection (), or has no DECam observation and non-detection by other shallower surveys. The remaining eight T dwarf candidates, all with clear DECam -band detections, comprise two newly identified wide binaries and six known wide systems (Gliese 570 A & D, K4V + T7.5, Burgasser et al. 2000; G 204-39 AB, M3+T6.5, Faherty et al. 2010; SDSS J1416+13 AB, sdL7+sdT7.5, Scholz 2010; Burningham et al. 2010; HIP 73786 AB, sdM4+sdT5.5, Murray et al. 2011; Zhang et al. 2019; LHS 6176 AB, K8V+T8p, Burningham et al. 2013; L 34-26 + COCONUTS-2b, M3V+T9, Zhang et al. 2021). Fig. 1 shows the PMs of these common PM pairs.
@@
The two new binaries, L 122-88 AB and UPM J10403551 AB have been independently reported recently. Marocco et al. (2024) classified UCAC3 52-1038 B (i.e. L 122-88 B) as a T6 dwarf based on photometric colours, and classified UCAC3 52-1038 A (i.e. L 122-88 A) as an M2 dwarf based on its low-resolution optical spectrum observed on 2021 December 25. Rothermich et al. (2024) classified UPM J1040–3551 A as an M3.8 dwarf based on photometric colours, and classified UPM J1040–3551 B as T7 dwarf based on a low resolution near infrared spectrum observed on 2023 June 30. Figure 2 shows the finding charts of these two new binaries. Their properties are listed in Table 1. The binary nature of these two newly identified pairs is corroborated by the concordance between the Gaia parallax-derived distances of the stellar components and the spectroscopically determined distances of their BD companions (Section 4.2).
@@
L 122-88 AB and UPM J10403551 AB are located at distances of 33 and 25 pc respectively. They were missed by earlier searches with WISE and AllWISE (Wright et al., 2010; Mainzer et al., 2011), possibly because T dwarf candidates within 20 pc were prioritised in those spectroscopic follow up programs, to achieve a more complete volume-limited sample (e.g., Kirkpatrick et al., 2011, 2021; Mace et al., 2013). Another seven known T dwarf binaries were selected as T dwarf candidates by Eqs 1–7, but were not recovered as wide binaries mainly because of large uncertainties of their CatWISE PMs. We recovered most of them by double the limits in Eqs 3 and 4, but no new binary T dwarf candidate was found.
Primary | L 122-88 A | UPM J10403551 A | Ref. |
@@
cProjected separation in Jacobi radius (, see Section 4.3.)
Name | SpT | UT date | Telescope | Instrument | Slit | Seeing | Airm | Wavelength | Resolution | |
(arcsec) | (arcsec) | (nm) | () | (sec) | ||||||
@@ 3 Spectroscopic observations
We conducted spectroscopic follow-up observations of L 122-88 AB and UPM J10403551 AB prior to their independent reporting. As spectra for L 122-88 B and UPM J10403551 A were not available in the literature, we continued our spectroscopic campaign to complete the analysis of these new binary systems. Table 2 summarises the details of our observations.
@@ 3.1 SOAR/Goodman
L 122-88 A and UPM J1040–3551 A were observed using the Goodman High Throughput Spectrograph (Goodman; Clemens et al., 2004) at the Southern Astrophysical Research Telescope (SOAR) in long-slit mode on 2023 December 29 (see Fig. 3). We used a slit and the 400 lines mm-1 grism in M2 mode, providing wavelength coverage from 492 to 897 nm with a resolving power of 2,000. Observations were conducted at airmasses of 1.89 and seeing of 1.55 for L 122-88 A, and airmasses of 1.04 and seeing of 1.45 for UPM J1040–3551 A.
Data reduction was performed using the Goodman HST Pipeline111https://soardocs.readthedocs.io/projects/goodman-pipeline/ for basic CCD reduction, including bias subtraction, flat-field correction, and cosmic ray removal. Further spectroscopic processing utilized tasks from the NOAO package in IRAF, including wavelength calibration, spectral extraction, telluric correction, and flux calibration. HIP 45167 served as the telluric standard star for flux calibration.
@@ 3.2 SOAR/TripleSpec
UPM J10403551 B and L 122-88 B were observed with the TripleSpec4.1 NIR Imaging Spectrograph (TripleSpec; Schlawin et al., 2014) at SOAR. TripleSpec provides simultaneous spectral coverage from 940 to 2470 nm (R3,500), spanning the entire photometric range (see Fig. 4). UPM J10403551 B was observed on 2024 March 26 (seeing: 091, airmass: 1.07), while L 122-88 B was observed on 2024 July 11 (seeing: 124, airmass: 1.35). We adopted the default ABBA dither pattern with eight 225-s exposures for UPM J1040–3551 B and twenty 200-s exposures for L 122-88 B.
Data reduction employed a modified version of the IDL-based Spextool pipeline (Cushing et al., 2004), adapted for SOAR. The process included flat-field correction, wavelength calibration using a CuHeAr arc lamp, removal of emission sky features by subtracting A and B exposures, and one-dimensional spectral extraction. Telluric correction and flux calibration were performed using observations of the standard star HIP54311. The resulting NIR spectra exhibit peak signal-to-noise ratios (SNR) of 10-15 at the centre of the -bands.
@@ 3.3 VLT/X-shooter
The X-shooter spectrograph (Vernet et al., 2011) at the Very Large Telescope comprises three spectroscopic arms (UVB, VIS, and NIR), each optimized for its spectral range (300-2500 nm). The optical to NIR spectrum of L 122-88 A was obtained using the X-shooter on 2024 June 16 (seeing: 15, airmass: 1.8). Figure 5 presents the combined, smoothed X-shooter spectrum of L 122-88 A.
L 122-88 B was observed with X-shooter on 2024 July 7 (seeing: 358) and 2024 July 13 (seeing: 129). Figure 6 shows the X-shooter spectrum of L 122-88 B, noting that the optical spectrum is contaminated by a nearby galaxy, in fact, we note the presence of a non-zero continuum in the 400-800 nm spectral region in which T dwarfs should barely have emit any flux (e.g., Burgasser et al., 2003b). The nearby galaxy is faint in the -band (see Fig. 2) but brighter than L 122-88 B in the optical. With a PM of 0368 yr-1 towards the north-east, L 122-88 B has moved to the top of the galaxy in 2024 when the X-shooter spectrum was taken. Meanwhile, the flux of the X-shooter spectrum beyond 2250 nm is poorly extracted/calibrated due to low SNR. These spectra were reduced using the ESO Reflex pipeline (Freudling et al., 2013).

### v3

# Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young [M4 + M4] + [T7 + T8] hierarchical quadruple

@@ Abstract
Benchmark brown dwarfs in wide binary systems are crucial for characterizing substellar objects and calibrating atmospheric and evolutionary models. However, brown dwarf benchmarks with subsolar metallicity, very cool temperatures, or suitability for dynamical mass measurements are rare, limiting our understanding across the full range of mass, age, and metallicity. We present the discovery of two new multiple systems containing T dwarf companions, identified through a targeted search using CatWISE2020 and Gaia catalogues. L 122-88 AB is a wide binary comprising a mildly metal-poor M2 dwarf and a T5 dwarf, separated by 215.6 arcsec at a distance of 33.1060.014 pc. Atmospheric model fitting to the near infrared spectrum of L 122-88 A suggests a mildly metal-poor composition ([Fe/H] = 0.2). UPM J10403551 AB is a candidate hierarchical quadruple system at 25.2830.013 pc, consisting of a likely astrometric binary of two M4 dwarfs and a probable unresolved spectral binary of T7 and T8 dwarfs, separated by 65.48 arcsec from the primary. The H emission detected in UPM J10403551 A indicates an age range of 0.3-2.0 Gyr. This age estimate suggests that the T8 component has a mass between 9 and 28 Jupiter masses, potentially classifying it as a planetary-mass object. These systems augment the sample of benchmark brown dwarfs, particularly in the underexplored regime of cool temperature, providing valuable opportunities for refining our understanding of substellar objects.
@@
planets and satellites: gaseous planets – binaries: general – brown dwarfs – stars: late-type – stars: low-mass1 Introduction
@@
Brown dwarfs (BDs) are the most recently discovered major population on the Hertzsprung-Russell diagram. They mostly form like stars but lack sufficient mass to sustain steady hydrogen fusion. Consequently, BDs represent the low-mass tail of the initial mass function and are crucial for constraining it (Kirkpatrick et al., 2024). Bridging the gap between stars and planets, BDs share similar physical and atmospheric properties with gas giant planets (Hatzes & Rauer, 2015), providing essential references for the characterization of exoplanets.
@@
To address the scientific objectives outlined above, samples of well-characterized BDs are required. However, BDs are challenging to characterize due to their faintness, cool temperatures, and the mass–age degeneracy inherent in their evolution. Without steady hydrogen burning, BDs cool over time, altering their spectral features and types (e.g. fig. 8 of Burrows et al., 2001). Therefore, benchmark BD samples with known mass, age, metallicity, and luminosity, spanning a wide parameter space, are needed to test atmospheric and evolutionary models. This need is especially pressing in the era of ongoing and upcoming deep surveys such as Euclid (Laureijs et al., 2011), the Roman Space Telescope (Spergel et al., 2015), the Rubin Observatory (LSST Science Collaboration et al., 2017), and the Chinese Space Station Telescope (CSST; Zhan, 2011).
@@
BDs in nearby open clusters or moving groups with well-established ages serve as valuable age benchmarks [e.g., Teide 1, (Rebolo et al., 1995) and Luhman 16 AB, (Luhman, 2013)]. However, these clusters and groups are predominantly young ( 1 Gyr) and relatively distant ( 100 pc). In contrast, field BDs constitute the majority of the BD population, and span a broader range of ages and metallicities, with many residing in the solar neighbourhood. Yet, direct measurements of age or mass for individual field BDs remain elusive. None the less, BDs that are wide companions to well-characterized stars provide ideal benchmarks [e.g., Gliese 229 B, (Nakajima et al., 1995) and Indi B, (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010)]. Binary systems are expected to have coeval formation, sharing common age, metallicity, and distance – parameters more readily determined for stars than for BDs. This characteristic makes such systems particularly valuable for BD studies.
@@
The T dwarf spectral class was established to classify BDs cool enough to exhibit methane absorption in their near-infrared (NIR) spectra (Burgasser et al., 2002; Geballe et al., 2002). T dwarfs have effective temperatures () between 500 and 1300 K (Kirkpatrick et al., 2021). They constitute the majority of field BDs (e.g. Zhang et al., 2019). However, T dwarfs are relatively rare in observations due to their cool temperatures and low luminosities. About 900 spectroscopically confirmed T dwarfs are known to date, mostly within 30 pc of the Sun (see summary in Best et al., 2024). Two-thirds of known T dwarfs have spectral types of T5–T9, owing to the rapid evolution of BDs through the L/T transition (e.g. Best et al., 2021).
@@
Currently, there are about 30 T dwarfs known with stellar companions (summarized in Best et al., 2024), with significant contributions from the Wide-field Infrared Survey Explorer (WISE; Wright et al., 2010). To expand the parameter space of benchmark T dwarfs and improve their characterization, we searched for late-type T dwarfs with wide stellar companions using the CatWISE2020 catalogue (Marocco et al., 2021) and the third data release (DR3) of Gaia (Gaia Collaboration et al., 2023). In this paper, we present the discovery of a wide binary of mildly metal-poor M2 + T5 dwarfs and a probable young hierarchical quadruple system of [M4 + M4] + [T7 + T8] dwarfs. Section 2 describes the selection of wide binary candidates. Spectroscopic observation and characterization of these two systems are presented in Sections 3 and 4, respectively. A summary and conclusions are presented in Section 5.
@@ 2 Selection of wide binaries
Wide binaries ( 100 au) exhibit negligible orbital motion over decadal time-scales due to their extremely long orbital periods ( 1 kyr). In the solar neighbourhood ( 100 pc), the proper motions (PMs) of wide binary components significantly exceed their orbital motions, resulting in apparent common PM over decades. This characteristic makes common PM an effective method for identifying nearby wide binaries (e.g., Zhang, 2019; Zhang et al., 2024).
T dwarfs emitting primarily in the infrared, only those in the solar neighbourhood are detectable by large-scale infrared surveys. The CatWISE catalogue, providing mid-infrared photometry and PMs for nearby objects, serves as a powerful tool for identifying nearby T dwarf wide binaries.
We focused on late-type T dwarfs, which exhibit redder colour (equation 1) and suffer less contamination, e.g. from poorly measured background objects, than earlier types. Candidate selection from CatWISE employed the following criteria:
In our selection, we required signal-to-noise ratio (SNR) of better than 5.0 for and detections (equation 2). SNR of 5.0 corresponds to an error of 0.22 in W1 or W2 magnitudes. To mitigate contamination from background stars in the low-resolution (2.75 arcsec pixel-1) unblurred coadds of the WISE images (unWISE; Lang, 2014) used for CatWISE, we excluded the Galactic plane (equation 5). We select objects without identified artefacts in both and band unWISE images (equation 7) and with only one blend component used in the fit (equation 6). Approximately 9,540 objects in CatWISE met these criteria (Eqs 1–7).
@@
Given the relatively large PM errors in CatWISE, we required significant PM (equation 3) but relaxed the PM constraint in the GCNS cross-match (equation 8). The distance limit of 40 pc (equation 9) corresponds to the -band detection limit of CatWISE for late T dwarfs (e.g. table 6 of Zhang et al., 2019). We limited the binary separation to within 10 arcmin (corresponding to 24,000 au at 40 pc), beyond which the number of random contaminations arises, partially due to the large PM errors in CatWISE.
@@
Our search yielded 15 candidate common PM pairs. After visual inspection using data from the Dark Energy Survey (Abbott et al., 2021) and the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Surveys (Dey et al., 2019), we excluded seven pairs of which the faint component has poor unWISE images, has Dark Energy Camera (DECam) deep -band non-detection (), or has no DECam observation and non-detection by other shallower surveys. The remaining eight T dwarf candidates, all with clear DECam -band detections, comprise two newly identified wide binaries and six known wide systems [Gliese 570 A & D, K4V + T7.5, (Burgasser et al., 2000); G 204-39 AB, M3+T6.5, (Faherty et al., 2010); SDSS J1416+13 AB, sdL7+sdT7.5, (Scholz, 2010; Burningham et al., 2010); HIP 73786 AB, sdM4+sdT5.5, (Murray et al., 2011; Zhang et al., 2019); LHS 6176 AB, K8V+T8p, (Burningham et al., 2013); L 34-26 + COCONUTS-2b, M3V+T9, (Zhang et al., 2021)]. Fig. 1 shows the PMs of these common PM pairs.
@@
The two new binaries, L 122-88 AB and UPM J10403551 AB have been independently reported recently. Marocco et al. (2024) classified UCAC3 52-1038 B (i.e. L 122-88 B) as a T6 dwarf based on photometric colours, and classified UCAC3 52-1038 A (i.e. L 122-88 A) as an M2 dwarf based on its low-resolution optical spectrum observed on 2021 December 25. Rothermich et al. (2024) classified UPM J1040–3551 A as an M3.8 dwarf based on photometric colours, and classified UPM J1040–3551 B as T7 dwarf based on a low-resolution NIR spectrum observed on 2023 June 30. Figure 2 shows the finding charts of these two new binaries. Their properties are listed in Table 1. The binary nature of these two newly identified pairs is corroborated by the concordance between the Gaia parallax-derived distances of the stellar components and the spectroscopically determined distances of their BD companions (Section 4.2).
@@
L 122-88 AB and UPM J10403551 AB are located at distances of 33 and 25 pc respectively. They were missed by earlier searches with WISE and AllWISE (Wright et al., 2010; Mainzer et al., 2011), possibly because T dwarf candidates within 20 pc were prioritized in those spectroscopic follow up programmes, to achieve a more complete volume-limited sample (e.g., Kirkpatrick et al., 2011, 2021; Mace et al., 2013). Another seven known T dwarf binaries were selected as T dwarf candidates by equations (1–7), but were not recovered as wide binaries mainly because of large uncertainties of their CatWISE PMs. We recovered most of them by double the limits in equations (3) and (4), but no new binary T dwarf candidate was found.
Primary | L 122-88 A | UPM J10403551 Aa/Ab | Ref. |
@@
cProjected separation in Jacobi radius (, see Section 4.3).
Name | SpT | UT date | Telescope | Instrument | Slit | Seeing | Airmass | Wavelength | Resolution | |
(arcsec) | (arcsec) | (nm) | () | (s) | ||||||
@@ 3 Spectroscopic observations
We conducted spectroscopic follow-up observations of L 122-88 AB and UPM J10403551 AB prior to their independent reporting. As spectra for L 122-88 B and UPM J10403551 A were not available in the literature, we continued our spectroscopic campaign to complete the analysis of these new binary systems. Table 2 summarizes the details of our observations.
@@ 3.1 SOAR/Goodman
L 122-88 A and UPM J1040–3551 A were observed using the Goodman High Throughput Spectrograph (Goodman; Clemens et al., 2004) at the Southern Astrophysical Research (SOAR) telescope in long-slit mode on 2023 December 29 (see Fig. 3). We used a slit and the 400 lines mm-1 grism in M2 mode, providing wavelength coverage from 492 to 897 nm with a resolving power of 2000. Observations were conducted at airmass of 1.89 and seeing of 1.55 for L 122-88 A, and airmass of 1.04 and seeing of 1.45 for UPM J1040–3551 A.
Data reduction was performed using the Goodman High Throughput Spectrograph Pipeline111https://soardocs.readthedocs.io/projects/goodman-pipeline/ for basic CCD reduction, including bias subtraction, flat-field correction, and cosmic ray removal. Further spectroscopic processing utilized tasks from the NOAO package in IRAF, including wavelength calibration, spectral extraction, telluric correction, and flux calibration. HIP 45167 served as the telluric standard star for flux calibration.
@@ 3.2 SOAR/TripleSpec
UPM J10403551 B and L 122-88 B were observed with the TripleSpec4.1 NIR Imaging Spectrograph (TripleSpec; Schlawin et al., 2014) at SOAR. TripleSpec provides simultaneous spectral coverage from 940 to 2470 nm (3500), spanning the entire photometric range (see Fig. 4). UPM J10403551 B was observed on 2024 March 26 (seeing: 0.91 arcsec, airmass: 1.07), while L 122-88 B was observed on 2024 July 11 (seeing: 1.24 arcsec, airmass: 1.35). We adopted the default ABBA dither pattern with eight 225-s exposures for UPM J1040–3551 B and twenty 200-s exposures for L 122-88 B.
Data reduction employed a modified version of the IDL-based Spextool pipeline (Cushing et al., 2004), adapted for SOAR. The process included flat-field correction, wavelength calibration using a CuHeAr arc lamp, removal of emission sky features by subtracting A and B exposures, and one-dimensional spectral extraction. Telluric correction and flux calibration were performed using observations of the standard star HIP 54311. The resulting NIR spectra exhibit peak SNRs of 10-15 at the centre of the bands.
@@ 3.3 VLT/X-shooter
The X-shooter spectrograph (Vernet et al., 2011) at the Very Large Telescope (VLT) comprises three spectroscopic arms [UV-Blue (UVB), Visible (VIS), and NIR], each optimized for its spectral range (300-2500 nm). The optical to NIR spectrum of L 122-88 A was obtained using the X-shooter on 2024 June 16 (seeing: 1.5 arcsec, airmass: 1.8). Figure 5 presents the combined, smoothed X-shooter spectrum of L 122-88 A.
L 122-88 B was observed with X-shooter on 2024 July 7 (seeing: 3.58 arcsec) and 2024 July 13 (seeing: 1.29 arcsec). Fig. 6 shows the X-shooter spectrum of L 122-88 B; noting that the optical spectrum is contaminated by a nearby galaxy, in fact, we note the presence of a non-zero continuum in the 400-800 nm spectral region in which T dwarfs should barely have emitted any flux (e.g., Burgasser et al., 2003b). The nearby galaxy is faint in the band (see Fig. 2) but brighter than L 122-88 B in the optical. With a PM of 0.368 arcsec yr-1 towards the north-east, L 122-88 B has moved to the top of the galaxy in 2024 when the X-shooter spectrum was taken. Meanwhile, the flux of the X-shooter spectrum beyond 2250 nm is poorly extracted/calibrated due to low SNR. These spectra were reduced using the European Southern Observatory Reflex pipeline (Freudling et al., 2013).